# Human-Centered Development of Guide Dog Robots: Quiet and Stable Locomotion Control

### v1

@@ V-B Hardware Experiments
For a more comprehensive evaluation, we conducted comparison experiments to benchmark our controllers with other methods. The first test involved forward walking over challenging terrains, including a whiteboard with soapy water, wooden blocks of varying dimensions, and a pile of cardboard boxes (see Fig. 6a). Across ten trials, our controller successfully navigated all terrains without a single failure, demonstrating strong robustness.
In the second experiment, we gradually increased the robot’s yaw velocity on a slippery board until it fell. Despite slipping during the tests (see Fig.6b), the robot successfully recovered and maintained balance at yaw angular velocities exceeding 3.0rad/s. To validate the significance of our MPC refinements, we compared our controller with state-of-the-art optimization-based controllers (Convex MPC [33] and Nonlinear MPC [34]) as well as a reinforcement learning (RL)-based method with a noise reduction reward, as described by Kim et al.[25]. For a fair comparison, the gait frequency of all controllers was set near 2Hz, except for the default controller due to limited low-level code access. The results show that, unlike other optimization-based controllers, ours maintained balance at significantly higher yaw rotational speeds. The achieved robustness is comparable to the RL method, which is known for its strong balance capabilities (see Fig. 6c).
In the third experiment, we measured acoustic noise while the robot performed straight walking on a flat surface, following the same procedure described in the paper (Section V.B). As shown in Fig. 6d, our method produced substantially quieter noise levels compared to all other methods.

### v2

@@ V-B Hardware Experiments
We conducted additional experiments to evaluate the generalization, robustness, and walking noise of our locomotion controller, comparing its performance against state-of-the-art optimization-based and reinforcement learning (RL) approaches.
V-B1 Challenging Terrain Walking
This experiment tested forward walking over diverse challenging terrains, including a whiteboard coated with soapy water, wooden blocks of varying dimensions, and a pile of cardboard boxes (see Fig. 6 (a)). Across ten trials, our controller successfully traversed all surfaces without a single failure, highlighting its strong robustness in unstructured environments.
V-B2 Turning on Slippery Surfaces
To evaluate the stability of our controller, we gradually increased the robot’s yaw velocity on a slippery board until failure occurred. Despite frequent slips (see Fig. 6 (b)), the robot maintained balance and successfully recovered at yaw angular velocities exceeding 3.0rad/s. We compared our controller against leading optimization-based methods - Convex MPC [33] and Nonlinear MPC [34] - as well as an RL-based method with a noise reduction reward, as described by Kim et al. [25]. For consistency, all controllers operated at a gait frequency of approximately 2Hz (except the default controller due to limited low-level code access). Our controller outperformed optimization-based methods in maximum stable yaw velocity and matched the robustness of the RL method (see Fig. 6 (c)).
V-B3 Acoustic Noise Evaluation
We quantified acoustic noise during straight-line walking on flat terrain, following the previous procedure. As illustrated in Fig. 6 (d), our controller consistently produced substantially lower noise levels compared to all baseline methods.